Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probablity ques

Expert replies
by nitug » Sat Aug 30, 2008 12:18 pm
hi all,

i almost always get probabilty ques wrong..plz do help me guys, till then i m looking and exploring the forums if someone has already pointed out the strategies to improve it..

Also solve this one ques for me,

The bowl contains chips of red and green color and no chips of any other color. If two chips are drawn from the basket at random without replacement, what is the probability that both chips will be green?
1. 20% of all chips in the basket are green
2. The ratio of the number of red chips to the number of green chips is 4:1


• Statement (1) ALONE is sufficient, but Statement (2) ALONE is not sufficient
• Statement (2) ALONE is sufficient, but Statement (1) ALONE is not sufficient
• BOTH statements TOGETHER are sufficient, but NEITHER statement ALONE is sufficient
• EACH statement ALONE is sufficient
• Statements (1) and (2) TOGETHER are NOT sufficient


i really appreciate this forum and its members!!

thanks,
nitug
Join the discussion
Source: — Data Sufficiency |

Re: probablity ques

by sudhir3127 » Sat Aug 30, 2008 1:01 pm
nitug wrote:hi all,

i almost always get probabilty ques wrong..plz do help me guys, till then i m looking and exploring the forums if someone has already pointed out the strategies to improve it..

Also solve this one ques for me,

The bowl contains chips of red and green color and no chips of any other color. If two chips are drawn from the basket at random without replacement, what is the probability that both chips will be green?
1. 20% of all chips in the basket are green
2. The ratio of the number of red chips to the number of green chips is 4:1


• Statement (1) ALONE is sufficient, but Statement (2) ALONE is not sufficient
• Statement (2) ALONE is sufficient, but Statement (1) ALONE is not sufficient
• BOTH statements TOGETHER are sufficient, but NEITHER statement ALONE is sufficient
• EACH statement ALONE is sufficient
• Statements (1) and (2) TOGETHER are NOT sufficient


i really appreciate this forum and its members!!

thanks,
nitug
IMO E
Join the discussion

by LSB » Sat Aug 30, 2008 1:34 pm
I think E as well. Both statements say the same thing. that 1 out of balls is green. However, this is a dependent prob question and we need to know how many balls are left once one is removed (or what the new ratios will be once one is removed)
Join the discussion

by sethids » Sat Aug 30, 2008 1:36 pm
Both the statements here give the same information.
20% chips are green, therefore 80% are red. Hence the ratio of red : green is 4 : 1.

Had it been a question of selection with replacement then the ratio would have been sufficient to answer. The probability of getting a green chip in one selection would have been 1/5. Because there are 2 selections, the probability of those 2 events happening simultaneously would have been 1/5 * 1/5 = 1/25.

But, here once we have made a selection (which still holds a probability of 1/5). We cannot determine the probability of the next selection unless we know the number of red and green chips or the modified ratio after one selection has been made.

Hence E.
Join the discussion

by nitug » Mon Sep 01, 2008 9:15 am
ans is E and again i made the mistake of not understanding the ques completely and missed the words "without replacement"
thanks for pointing that out sethids.
Join the discussion